# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v2

@@ F.2 Four-point functions
where $\left\langle\left\langle A,B\right\rangle\right\rangle_{\star}=\left\langle AB\right\rangle_{\star}-\left\langle A\right\rangle_{\star}\left\langle B\right\rangle_{\star}$, $I(t_{1},t_{2},t_{3},t_{4})=\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})$. Expanding the action to second order around the saddle point, the path integral becomes
@@
In principle, we have all the necessary components to evaluate this expression: the single-site dynamics are known via the saddle-point condition, and the Hessian blocks are given in terms of connected correlation functions in this single-site process. We could then take the temporal separation limits of Eq. 34 to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$. The challenge is that each Hessian block is a complicated time2-by-time2 matrix that, for general values of the time variables, has no analytic inverse. To circumvent this problem, we show in Appendix G that the temporal limits of Eq. 34 can be taken before taking the inverses. Under these temporal limits, the Hessian blocks are
@@
$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(\tau_{1})S^{\phi}_{\star}(\tau_{2}),$
@@
where we used the fact that multiplication by $i\hat{x}(t)$ is equivalent to a functional derivative with respect to a source at time $t$. Since the relevant quantities are now translation-invariant, i.e., depend only on $\tau_{1}$ and $\tau_{2}$, we can transform to Fourier space. In this representation, each Hessian block becomes a frequency-dependent scalar,
@@
In summary, the full Hessian can be replaced by a two-by-two, frequency-dependent Hesssian given by
@@
where we simplified notation by suppressing frequency arguments via the shorthand $C^{\phi}_{12}=C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2})$ and $S^{\phi}_{12}=S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})$. The Fourier-space function $\Psi_{\star}^{\phi}(\bm{\omega})$ is the upper-left element of the inverse of this matrix. Performing the two-by-two matrix inversion gives
@@
which agrees with the two-site cavity result (Eq. 72).
@@ F.3 Adding sources
Our goal is to compute
@@
$Z[\mathcal{J}]=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp\left(-N\hat{U}U+N\hat{U}\mathcal{J}-NS[\hat{C};U]\right).$ (91)
Instead of taking the second derivative with respect to $\mathcal{J}$, we set $\mathcal{J}=0$ and compute the fluctuations of $\hat{U}$ in the augmented theory whose path integral is
@@
The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point.
Consider a general intensive action $\mathcal{S}[\bm{C};\mathcal{J}]$ where $\bm{C}$ is a collection of fields (e.g., $\bm{C}=\{C^{\phi},\hat{C}^{\phi}\}$), and $\mathcal{J}$ is a source. Using a Schur complement to compute this element, we obtain
@@
with all quantities evaluated at the saddle point. By applying this formula to the action of the i.i.d. model with the source term $\frac{1}{2}x\mathcal{J}x$ added to the single-site path integral (Eq. 76), we recover the expression for $\Psi^{x}_{\star}(\bm{\tau})$ from the two-site cavity method (Clark et al. [30]; Eq. 73). For actions of the form $\mathcal{S}[\bm{C};\mathcal{J}]=-\mathcal{J}C_{a}+\text{$\mathcal{J}$-independent terms}$, this formula reduces to $({\delta^{2}}/{\delta\mathcal{J}^{2}})\log Z|_{\mathcal{J}=0}=[\bm{H}^{-1}]_{aa},$ as expected.
We now apply Eq. 96 to the action of the i.i.d. connectivity model with a source term for correlations among pre-activations. The action is
@@
Computing the necessary quantities and taking the temporal limits of Eq. 34, we obtain, in Fourier space,
@@
where $C^{x\phi}_{12}=C_{\star}^{x\phi}(\omega_{1})C_{\star}^{x\phi}(\omega_{2})$ and $S^{x}_{12}=S^{x}_{\star}(\omega_{1})S^{x}_{\star}(\omega_{2})$. Substituting these into Eq. 96 and using the frequency-dependent Hessian from Eq. 88, we obtain
@@
in agreement with the two-site cavity result of Clark et al. [30], given in Eq. 73.
@@ F.4 Other kinds of path integrals
Different path-integral formalisms have been applied to neural network models depending on the underlying dynamics. The Doi-Peliti formalism [93] derives path integrals from master equations for systems with non-negative integer degrees of freedom (such as spiking neural networks) that evolve according to master equations, using an operator formalism as an intermediate step. In contrast, we use the MSRDJ formalism for continuous-time stochastic differential equations, which naturally describes our rate-based network dynamics. For a comprehensive discussion of different path-integral approaches in neural network theory, see [60].
@@ Appendix G Temporal limits and inverses
Here, we demonstrate, for the i.i.d. model that the temporal limits in Eq. 34 can be taken before the inverses. This allows us to work with a frequency-dependent Hessian. In this section, we use Einstein notation for integrals over time variables (i.e., repeated time indices are integrated over). We would like to evaluate
$\Psi(\tau_{1},\tau_{2})=U^{-1}(t_{1},t_{2},s_{1},s_{2})K(s_{1},s_{2},s_{3},s_{4})U^{-1}(t_{1}+\tau_{1},t_{2}+\tau_{2},s_{3},s_{4})$ (100)
in the limits Eq. 34, where $U(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ and $K(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ (note that $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0$ due to the vanishing of correlation functions involving only the conjugate field). Restating Eq. 81, we have
@@
We start by examining the properties of $(t_{1},t_{2},t_{3},t_{4}). UU$ vanishes if one of the four time points is far (measured on the timescale of $C^{\phi}_{\star}(\tau)$) from the other three. For $t_{1}$ or $t_{2}$, this is a consequence of the perturbation associated with $\hat{x}(t_{1})$ or $\hat{x}(t_{2})$ being far away; for $t_{3}$ or $t_{4}$, this is a consequence of $\langle\phi(t_{i})\rangle_{\star}=0$ separating from the expectation. $U$ also vanishes if there is more than one time point which is far from the remaining ones by the same arguments. Thus, the time points need to be close in pairs. However, if $t_{1},t_{2}$ are close and $t_{3},t_{4}$ are close, but the pairs are distant, $U$ still vanishes since the perturbations associated with $\hat{x}(t_{i})$ are both far. This leaves three possibilities for a non-vanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
Separating the three possibilities into two subdomains where all time points are close ($c$) and where the time points belong to either of the two pairwise separate ($s$) possibilities, we write $U=U_{s}+U_{c}$ where $U_{s}$ and $U_{c}$ are defined to vanish on the respective other domain. On the separated domain, $U$ simplifies to
@@
where we used that the two distant contributions are non-overlapping to add them.
@@
For Eq. 100, we need $U^{-1}_{s}$ due to the limits Eq. 34. The fact that $U$ reduces to $U_{s}$ in the limits Eq. 34 justifies taking the limits before inverting $U$. Furthermore, these limits imply that, in Eq. 100, $s_{1}$ and $s_{2}$, as well as $s_{3}$ and $s_{4}$, are far apart. In this case, $K$ simplifies to
@@
This justifies taking the limits Eq. 34 for $K$.
@@ Appendix H Partially symmetric disorder
$\left\langle J_{ij}J_{k\ell}\right\rangle=g^{2}\delta_{ik}\delta_{j\ell}+g^{2}\rho\delta_{i\ell}\delta_{jk}.$ (105)
Here, $\rho\in[-1,1]$ is a parameter that controls the degree of symmetry in the connectivity. When $\rho=1$, the connectivity is fully symmetric, while $\rho=-1$ corresponds to fully antisymmetric connectivity. The case $\rho=0$ recovers the i.i.d. model. We express this partially symmetric connectivity as a linear combination of two matrices, $\bm{X}$ and $\bm{Y}$, that are i.i.d. random with mean zero and variance $1/N$:
$J_{ij}=\sqrt{{g^{2}-|\sigma|}}X_{ij}+\sqrt{\frac{|\sigma|}{2}}(Y_{ij}+\text{sgn}(\sigma)Y_{ji}),$ (106)
@@
$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg{(}i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{ij}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{ij}Y_{ij}(\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j})\bigg{)}.$ (107)
@@ H.1 Two-point functions
Setting these derivatives to zero yields the saddle-point conditions:
@@
$T[x](t)$ $=\eta^{x}(t)+\sigma\left[S^{\phi}_{\star}\ast\phi\right](t),$ (120)
@@
where $\ast$ denotes convolution. The symmetric structure provides a convolutional, nonlinear self-coupling $\sigma[S^{\phi}_{\star}\ast\phi](t)$ in the single-site dynamics. The two-point correlation and response functions, $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$, must be determined self-consistently within this single site picture.
@@ H.2 Four-point functions
Inverting this matrix and isolating the upper-left element, we obtain, in agreement with Clark et al. [30],
@@ Appendix I Path-integral calculation of two- and four-point functions for the random-mode model
and their conjugates $\hat{z}_{a}(t)$. We use indices $i,j$ for neurons and $a,b$ for latent variables. The introduction of these latent variables makes the action linear in $\bm{L}$ and $\bm{R}$:
@@
We can now easily integrate out $\bm{L}$ and $\bm{R}$. To factorize the action over indices spanning extensive dimensions, we introduce the field $C^{\phi}(t_{1},t_{2})$ (Eq. 17) with conjugate $\hat{Q}(t_{1},t_{2})$ as well as
@@ I.1 Two-point functions
showing that the neuronal and latent-variable single-site processes have mutually referential statistics  [^8] Note that the variance of $\eta^{z}(t)$ is scaled by $D^{2}$; equivalently, we could factor out $D^{2}$ from the variance and include $D$ as a multiplicative factor in the latent-variable single-site dynamics (Eq. 140).
Solving for $Q_{\star}(t_{1},t_{2})$ by combining Eq. 138, Eq. 140, and Eq. 142, we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq. 10. Consolidating these results yields Eq. 36c.
@@ I.2 Four-point functions
Figure 10: Analysis of dimension of activity and covariance spectrum in the low-effective-rank limit. (a) Top: Dimension of activity $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ ($\alpha=1$) for various coupling strengths $g_{\text{eff}}$. Component strengths $D_{a}$ are constant for $a/N\leq\text{PR}^{D}$ and 0 otherwise. Small dots: individual simulations; large dots: means over 10 simulations; solid lines: theoretical predictions; dotted lines: small-$\text{PR}^{D}$ expansion, $K^{\phi}(g_{\text{eff}})\mathrm{PR}^{D}$. Inset: Ratio of proportionality factor $K^{\phi}(g_{\text{eff}})$ to $\text{PR}^{\phi}_{\text{i.i.d.}}$ versus $g_{\text{eff}}$. Bottom: Same analysis for pre-activations. (b) Top: Rank-ordered eigenvalue plots for activation covariance matrix $C^{\phi}_{ij}(0)$. Dotted vertical lines indicate $N\,\text{PR}^{D}$. Bottom: Same for pre-activations. All simulations use $N=6500$ neurons.
@@
The upper-left block, corresponding to the latent variables, is related to the Hessian of an i.i.d. network with linear dynamics and a heterogeneous distribution of single-neuron scale factors $D_{a}$; the lower-right block, corresponding to the neurons, is related to the Hessian of an i.i.d. network of nonlinear neurons (Eq. 88). They are coupled through the simple off-diagonal blocks, reflecting the interaction between neurons and latent variables.
This yields the expression for $\Psi^{\phi}_{\star}$ (Eq. 38). To obtain $\Psi^{x}_{\star}$, we add $\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral (Eq. 130) and compute
@@
Inserting these into Eq. 96 we obtain
@@
where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ (Eq. 73b with $g=g_{\text{eff}}$). Here, we have used the following formula to invert the Hessian:
@@
where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}(x_{i}(t))$. The resulting action and single-site path integrals for this system are
@@
The self-consistency condition Eq. 155c is identical to that of an i.i.d. network, but with an average over $\bm{\theta}\sim P(\bm{\theta})$, in addition to the usual Gaussian average over $\eta^{x}(t)$, to account for single-neuron heterogeneity.
@@
To write down the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$:
@@
where the new $\mathfrak{F}$raktur variables are defined in Eq. 156. We invert the Hessian using Eq. 146 and identify the upper-left element as $\Phi_{\bm{\theta}_{i}}(\bm{\omega})$:
@@
we obtain for the normalized variables $\phi_{i}(t)$:
@@
Upon specializing the above general results to the case where $\bm{\theta}_{i}$ contains a single gain parameter, $G_{i}$, we obtain the four-point functions for unnormalized (Eq. 46a) and normalized (Eq. 46b) variables given in the main text.
@@
For “random readout” gains with the same distribution as the recurrent gains (see main text), the four-point function is
@@
We conclude with a note on gain modulation in recurrent neural networks. It has been shown that modulating single-neuron gains while keeping synaptic weights fixed has great expressive power over the dynamics of recurrent neural networks [95]. Here, we consider random gains rather than judiciously chosen or learned gains. In particular, they are sampled independently of the connectivity. However, our framework could be extended to model gains that are related to the connectivity and, more ambitiously, could be extended to model learned gain parameters (or, as explored in the Discussion, learned connectivity).
@@
To generate matrices $\bm{L}$ and $\bm{R}$ with appropriate correlation structure, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$:
@@
We begin with the path integral containing latent variables, insert this parameterization of $\bm{L}$ and $\bm{R}$, then and integrate out $\bm{X}_{1}$, $\bm{X}_{2}$, and $\bm{Y}$. To factorize the action over indices spanning extensive dimensions, we introduce
@@
The resulting path integral, action, and single-site path integrals are given by:
@@
We first calculate the saddle point conditions by taking the derivatives of the action:
@@
The saddle-point conditions yield:

### v3

@@ F.2 Four-point functions
where $\left\langle\left\langle A,B\right\rangle\right\rangle_{\star}=\left\langle AB\right\rangle_{\star}-\left\langle A\right\rangle_{\star}\left\langle B\right\rangle_{\star}$ and $I(t_{1},t_{2},t_{3},t_{4})=\delta(t_{1}-t_{3})\delta(t_{2}-t_{4})+\delta(t_{1}-t_{4})\delta(t_{2}-t_{3})$. Expanding the action to second order around the saddle point, the path integral becomes
@@
In principle, we have all the necessary components to evaluate this expression: The single-site dynamics are known via the saddle-point condition, and the Hessian blocks are given in terms of connected correlation functions in this single-site process. We could then take the temporal separation limits of Eq. (35) to obtain $\Psi^{\phi}_{\star}(\bm{\tau})$. The challenge is that each Hessian block is a complicated time2-by-time2 matrix that, for general values of the time variables, has no analytic inverse. To circumvent this problem, we show in Appendix G that the temporal limits of Eq. (35) can be taken before taking the inverses. Under these temporal limits, the Hessian blocks are
@@
$H_{C^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})=-\delta(\tau_{1})\delta(\tau_{2})+g^{2}S^{\phi}_{\star}(\tau_{1})S^{\phi}_{\star}(\tau_{2}),$ (85)
@@
where we used the fact that multiplication by $i\hat{x}(t)$ is equivalent to a functional derivative with respect to a source at time $t$. Since the relevant quantities are now translation invariant, i.e., depend only on $\tau_{1}$ and $\tau_{2}$, we can transform to Fourier space. In this representation, each Hessian block becomes a frequency-dependent scalar,
@@
In summary, the full Hessian can be replaced by a $2\times 2$, frequency-dependent Hessian given by
@@
where we simplified notation by suppressing frequency arguments via the shorthand $C^{\phi}_{12}=C^{\phi}_{\star}(\omega_{1})C^{\phi}_{\star}(\omega_{2})$ and $S^{\phi}_{12}=S^{\phi}_{\star}(\omega_{1})S^{\phi}_{\star}(\omega_{2})$. The Fourier-space function $\Psi_{\star}^{\phi}(\bm{\omega})$ is the upper-left element of the inverse of this matrix. Performing the $2\times 2$ matrix inversion gives
@@
which agrees with the two-site cavity result [Eq. (73)].
@@ F.3 Adding sources
Consider a general intensive action $\mathcal{S}[\bm{C};\mathcal{J}]$, where $\bm{C}$ is a collection of fields (e.g., $\bm{C}=\{C^{\phi},\hat{C}^{\phi}\}$), and $\mathcal{J}$ is a source. Our goal is to compute
@@
$Z[\mathcal{J}]=\int\mathcal{D}\bm{C}\int\mathcal{D}U\int\mathcal{D}\hat{U}\exp\left(-N\hat{U}U+N\hat{U}\mathcal{J}-NS[\bm{C};U]\right).$ (93)
Instead of taking the second derivative with respect to $\mathcal{J}$, we set $\mathcal{J}=0$ and compute the fluctuations of $\hat{U}$, measured by Eq. (92), in the augmented theory whose path integral is
@@
The fluctuations of $\hat{U}$ are given by the bottom-right element of the inverse Hessian evaluated at the saddle point. Using a Schur complement to compute this element, we obtain
@@
with all quantities evaluated at the saddle point. By applying this formula to the action of the i.i.d. model with the source term $\frac{1}{2}x\mathcal{J}x$ added to the single-site path integral [Eq. (77)], we recover the expression for $\Psi^{x}_{\star}(\bm{\tau})$ from the two-site cavity method [Clark et al. [39]; Eq. (74)]. For actions of the form $\mathcal{S}[\bm{C};\mathcal{J}]=-\mathcal{J}C_{a}+\text{$\mathcal{J}$-independent terms}$, this formula reduces to $({\delta^{2}}/{\delta\mathcal{J}^{2}})\log Z|_{\mathcal{J}=0}=[\bm{H}^{-1}]_{aa},$ as expected.
We now apply Eq. (98) to the action of the i.i.d. connectivity model with a source term for correlations among preactivations. The action is
@@
Computing the necessary quantities and taking the temporal limits of Eq. (35), we obtain, in Fourier space,
@@
where $C^{x\phi}_{12}=C_{\star}^{x\phi}(\omega_{1})C_{\star}^{x\phi}(\omega_{2})$ and $S^{x}_{12}=S^{x}_{\star}(\omega_{1})S^{x}_{\star}(\omega_{2})$. Substituting these into Eq. (98) and using the frequency-dependent Hessian from Eq. (90), we obtain
@@
in agreement with the two-site cavity result of Clark et al. (Ref. [39]), given in Eq. (74).
@@ F.4 Other kinds of path integrals
Different path-integral formalisms have been applied to neural-network models depending on the underlying dynamics. The Doi-Peliti formalism [100] derives path integrals from master equations for systems with non-negative integer degrees of freedom (such as spiking neural networks) that evolve according to master equations, using an operator formalism as an intermediate step. In contrast, we use the MSRJD formalism for continuous-time stochastic differential equations, which naturally describes our rate-based network dynamics. For a comprehensive discussion of different path-integral approaches in neural-network theory, see Ref. [69].
@@ Appendix G Temporal limits and inverses
Here, we demonstrate, for the i.i.d. model, that the temporal limits in Eq. (35) can be taken before the inverses. This allows us to work with a frequency-dependent Hessian. In this section, we use Einstein notation for integrals over time variables (i.e., repeated time indices are integrated over). We would like to evaluate
$\Psi^{\phi}(\tau_{1},\tau_{2})=U^{-1}(t_{1},t_{2},s_{1},s_{2})K(s_{1},s_{2},s_{3},s_{4})U^{-1}(t_{1}+\tau_{1},t_{2}+\tau_{2},s_{3},s_{4})$ (102)
in the limits of Eq. (35), where $U(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ and $K(t_{1},t_{2},t_{3},t_{4})=H_{\hat{C}^{\phi}\hat{C}^{\phi}}(t_{1},t_{2},t_{3},t_{4})$ [note that $H_{C^{\phi}C^{\phi}}(t_{1},t_{2},t_{3},t_{4})=0$ due to the vanishing of correlation functions involving only the conjugate field]. Restating Eq. (82), we have
@@
We start by examining the properties of $(t_{1},t_{2},t_{3},t_{4}). UU$ vanishes if one of the four time points is far [measured on the timescale of $C^{\phi}_{\star}(\tau)$] from the other three. For $t_{1}$ or $t_{2}$, this is a consequence of the perturbation associated with $\hat{x}(t_{1})$ or $\hat{x}(t_{2})$ being far away; for $t_{3}$ or $t_{4}$, this is a consequence of $\langle\phi(t_{i})\rangle_{\star}=0$ separating from the expectation. $U$ also vanishes if there is more than one time point which is far from the remaining ones by the same arguments. Thus, the time points need to be close in pairs. However, if $t_{1},t_{2}$ are close and $t_{3},t_{4}$ are close, but the pairs are distant, $U$ still vanishes since the perturbations associated with $\hat{x}(t_{i})$ are both far. This leaves three possibilities for a nonvanishing $U$: $t_{1},t_{3}$ close and $t_{2},t_{4}$ close, but the pairs are distant; $t_{1},t_{4}$ close and $t_{2},t_{3}$ close, but the pairs are distant; or all time points are close.
Separating the three possibilities into two subdomains where all time points are close ($c$) and where the time points belong to either of the two pairwise separate ($s$) possibilities, we write $U=U_{s}+U_{c}$, where $U_{s}$ and $U_{c}$ are defined to vanish on the respective other domain. On the separated domain, $U$ simplifies to
@@
where we used that the two distant contributions are nonoverlapping to add them.
@@
For Eq. (102), we need $U^{-1}_{s}$ due to the limits of Eq. (35). The fact that $U$ reduces to $U_{s}$ in the limits of Eq. (35) justifies taking the limits before inverting $U$. Furthermore, these limits imply that, in Eq. (102), $s_{1}$ and $s_{2}$, as well as $s_{3}$ and $s_{4}$, are far apart. In this case, $K$ simplifies to
@@
This justifies taking the limits of Eq. (35) for $K$.
@@ Appendix H Partially symmetric disorder
$\left\langle J_{ij}J_{k\ell}\right\rangle=\frac{g^{2}}{N}\delta_{ik}\delta_{j\ell}+\frac{g^{2}\rho}{N}\delta_{i\ell}\delta_{jk}.$ (107)
Here, $\rho\in[-1,1]$ is a parameter that controls the degree of symmetry in the connectivity. When $\rho=1$, the connectivity is fully symmetric, while $\rho=-1$ corresponds to fully antisymmetric connectivity. The case $\rho=0$ recovers the i.i.d. model. We express this partially symmetric connectivity as a linear combination of two matrices, $\bm{X}$ and $\bm{Y}$, that are i.i.d. random with mean zero and variance $1/N$,
$J_{ij}=\sqrt{{g^{2}-|\sigma|}}X_{ij}+\sqrt{\frac{|\sigma|}{2}}[Y_{ij}+\text{sgn}(\sigma)Y_{ji}],$ (108)
@@
$Z[\bm{X},\bm{Y}]=\int\mathcal{D}\bm{x}\int\mathcal{D}\hat{\bm{x}}\exp\bigg(i\sum_{i=1}^{N}\hat{x}_{i}T[x_{i}]-i\sqrt{{g^{2}-|\sigma|}}\sum_{i,j=1}^{N}X_{ij}\hat{x}_{i}\phi_{j}-i\sqrt{\frac{|\sigma|}{2}}\sum_{i,j=1}^{N}Y_{ij}[\hat{x}_{i}\phi_{j}+\text{sgn}(\sigma)\phi_{i}\hat{x}_{j}]\bigg).$ (109)
@@ H.1 Two-point functions
Setting these derivatives to zero yields the saddle-point conditions,
@@
$T[x](t)$ $=\eta^{x}(t)+\sigma\left[S^{\phi}_{\star}\circ\phi\right](t),$ (122)
@@
where $\circ$ denotes convolution. The symmetric structure provides a convolutional, nonlinear self-coupling $\sigma[S^{\phi}_{\star}\circ\phi](t)$ in the single-site dynamics. The two-point correlation and response functions $C^{\phi}_{\star}(\tau)$ and $S^{\phi}_{\star}(\tau)$ must be determined self-consistently within this single-site picture.
@@ H.2 Four-point functions
Inverting this matrix and isolating the upper-left element, we obtain, in agreement with Clark et al. (Ref. [39]),
@@ Appendix I Path-integral calculation of two- and four-point functions for the random-mode model
and their conjugates $\hat{z}_{a}(t)$. We use indices $i,j$ for neurons and $a,b$ for latent variables. The introduction of these latent variables makes the action linear in $\bm{L}$ and $\bm{R}$,
@@
We can now easily integrate out $\bm{L}$ and $\bm{R}$. To factorize the action over indices spanning extensive dimensions, we introduce the field $C^{\phi}(t_{1},t_{2})$ [Eq. (17)] with conjugate $\hat{Q}(t_{1},t_{2})$ as well as
@@ I.1 Two-point functions
showing that the neuronal and latent-variable single-site processes have mutually referential statistics  [^8] Note that the variance of $\eta^{z}(t)$ is scaled by $D^{2}$; equivalently, we could factor out $D^{2}$ from the variance and include $D$ as a multiplicative factor in the latent-variable single-site dynamics [Eq. (142)].
Solving for $Q_{\star}(t_{1},t_{2})$ by combining Eqs. (140), (142), and (144), we obtain $Q_{\star}(t_{1},t_{2})=\alpha r_{2}C^{\phi}_{\star}(t_{1},t_{2}),$ with $r_{n}$ defined by Eq. (10). Consolidating these results yields Eq. (37c).
@@ I.2 Four-point functions
Figure 10: Analysis of dimension of activity and covariance spectrum in the low-effective-rank limit. (a) Top: dimension of activity $\text{PR}^{\phi}$ versus effective rank $\alpha\text{PR}^{D}$ ($\alpha=1$) for various coupling strengths $g_{\text{eff}}$. Component strengths $D_{a}$ are constant for $a/N\leq\text{PR}^{D}$ and 0 otherwise. Small dots, individual simulations; large dots: means over ten simulations; solid lines, theoretical predictions; dotted lines, small-$\text{PR}^{D}$ expansion, $K^{\phi}(g_{\text{eff}})\mathrm{PR}^{D}$. Inset: ratio of proportionality factor $K^{\phi}(g_{\text{eff}})$ to $\text{PR}^{\phi}_{\text{i.i.d.}}$ versus $g_{\text{eff}}$. Bottom: same analysis for preactivations. (b) Top: rank-ordered eigenvalue plots for activation covariance matrix $C^{\phi}_{ij}(0)$. Dotted vertical lines indicate $N\,\text{PR}^{D}$. Bottom: same for preactivations. All simulations use $N=6500$ neurons.
@@
The upper-left block, corresponding to the latent variables, is related to the Hessian of an i.i.d. network with linear dynamics and a heterogeneous distribution of single-neuron scale factors $D_{a}$; the lower-right block, corresponding to the neurons, is related to the Hessian of an i.i.d. network of nonlinear neurons [Eq. (90)]. They are coupled through the simple off diagonal blocks, reflecting the interaction between neurons and latent variables.
This yields the expression for $\Psi^{\phi}_{\star}$ [Eq. (39)]. To obtain $\Psi^{x}_{\star}$, we add $\frac{1}{2}x\mathcal{J}x$ to the neuronal single-site path integral [Eq. (132)] and compute
@@
Inserting these into Eq. (98) we obtain
@@
where $U={g_{\text{eff}}^{2}S^{x}_{12}}/({1-g_{\text{eff}}^{2}S^{\phi}_{12}})$ [Eq. (74b) with $g=g_{\text{eff}}$]. Here, we have used the following formula to invert the Hessian:
@@
where $\Phi_{\bm{\theta}_{i}}(t)=\Phi_{\bm{\theta}_{i}}\bm{(}x_{i}(t)\bm{)}$. The resulting action and single-site path integrals for this system are
@@
The self-consistency condition Eq. (157c) is identical to that of an i.i.d. network, but with an average over $\bm{\theta}\sim P(\bm{\theta})$, in addition to the usual Gaussian average over $\eta^{x}(t)$, to account for single-neuron heterogeneity.
@@
To write down the solution for the four-point function, we define new two-frequency correlation functions with an outer average over $\bm{\theta}\sim P(\bm{\theta})$,
@@
where the new Fraktur variables are defined in Eq. (158). We invert the Hessian using Eq. (148) and identify the upper-left element as $\Phi_{\bm{\theta}_{i}}(\bm{\omega})$,
@@
we obtain for the normalized variables $\phi_{i}(t)$,
@@
Upon specializing the above general results to the case where $\bm{\theta}_{i}$ contains a single gain parameter $G_{i}$, we obtain the four-point functions for unnormalized [Eq. (47a)] and normalized [Eq. (47b)] variables given in the main text.
@@
For “random-readout” gains with the same distribution as the recurrent gains (see main text), the four-point function is
@@
We conclude with a note on gain modulation in recurrent neural networks. It has been shown that modulating single-neuron gains while keeping synaptic weights fixed has great expressive power over the dynamics of recurrent neural networks [102]. Here, we consider random gains rather than judiciously chosen or learned gains. In particular, they are sampled independent of the connectivity. However, our framework could be extended to model gains that are related to the connectivity and, more ambitiously, could be extended to model learned gain parameters (or, as explored in the Discussion, learned connectivity).
@@
To generate matrices $\bm{L}$ and $\bm{R}$ with appropriate correlation structure, we express them in terms of independent Gaussian random matrices $\bm{X}^{1},\bm{X}^{2},$ and $\bm{Y}$, each with variance $1/N$,
@@
We begin with the path integral containing latent variables, insert this parametrization of $\bm{L}$ and $\bm{R}$, and then integrate out $\bm{X}_{1}$, $\bm{X}_{2}$, and $\bm{Y}$. To factorize the action over indices spanning extensive dimensions, we introduce
@@
The resulting path integral, action, and single-site path integrals are given by
@@
We first calculate the saddle-point conditions by taking the derivatives of the action,
@@
The saddle-point conditions yield